OpenStudy (anonymous):

The volume of a hexagonal prism is 10square root 3 in3. What is the height of the prism if the base area is 5square root 3 in3?

OpenStudy (anonymous):

\[h = \frac{10\sqrt{3}}{5\sqrt{3}}\]

OpenStudy (anonymous):

Can you solve that?

OpenStudy (anonymous):

Yeah